# Total Failure of Approachability at Successors of Singulars of Countable Cofinality

### v1

@@ Abstract.
The notion of independence is fundamental to set-theoretic research. It refers to the phenomenon that many set-theoretic statements are neither proven nor refuted by $\mathsf{ZFC}$, the most commonly accepted axioms of set theory. One particular theme in the study of independence statements is the fact that successors of singular cardinals are often much less malleable than successors of regular cardinals – recall that a cardinal $\kappa$ is singular if we can write $\kappa=\bigcup_{i\in x}y_{i}$, where $x$ and every $y_{i}$ has size $\kappa$ and regular otherwise. Additionally, many independence results cannot be transferred from regular cardinals to singular cardinals. The most famous example of this phenomenon is the combination of two results due to William Easton and Saharon Shelah: Easton showed that the continuum function – mapping a cardinal $\kappa$ to the size of its powerset – when restricted to the class of regular cardinals can be equal to any function, provided that function is nondecreasing and maps every cardinal to a cardinal with a larger cofinality. On the other hand, Shelah showed that, assuming $2^{\aleph_{n}}<\aleph_{\omega}$ for every $n<\omega$, $2^{\aleph_{\omega}}<\min(\aleph_{(2^{\aleph_{0}})^{+}},\aleph_{\omega_{4}})$, so Easton’s result cannot be replicated for the class of all cardinals.
@@
Another example highlighting the difference between successors of singular cardinals and successors of regular cardinals is Shelah’s approachability ideal: Shelah introduced the ideal $I[\kappa^{+}]$ in [She79] in order to study the indestructibility of stationary subsets of $\kappa^{+}$ both for singular and regular cardinals $\kappa$. Despite the definition of $I[\kappa^{+}]$ being the same in both cases, its behavior differs greatly depending on whether $\kappa$ is regular or singular: When $\kappa$ is a regular cardinal, it is commonly referred to as the critial cofinality, since the set $E_{<\kappa}^{\kappa^{+}}$ (consisting of all ordinals below $\kappa^{+}$ with cofinality $\kappa$) is always a member of $I[\kappa^{+}]$ and so the behavior of $I[\kappa^{+}]$ is determined by which subsets of $E_{\kappa}^{\kappa^{+}}$ it contains. On the other hand, if $\kappa$ is singular, it was only proven that the set $E_{\leq\operatorname{cf}(\kappa)}^{\kappa^{+}}$ belongs to $I[\kappa^{+}]$ and so many more options regarding the behavior of $I[\kappa^{+}]$ seem possible.
@@
Recently, in joint work of Maxwell Levine and the author (see [JL25]), we solved an almost four-decades-old open problem due to Saharon Shelah by constructing a model – for an arbitrary $n\in(0,\omega)$ – where $E_{\aleph_{n}}^{\aleph_{\omega+1}}\notin I[\aleph_{\omega+1}]$. Previously it had only been known that there could be stationarily many non-approachable points of cofinality $\aleph_{1}$ in $\aleph_{\omega+1}$. Our result begs the natural question (which was also raised by Mitchell in [Mit09] and by Dobrinen in private communication) whether it is consistent that $E_{\mu}^{\kappa^{+}}\notin I[\kappa^{+}]$ for every $\mu\in(\aleph_{0},\kappa)$ (where $\kappa$ is singular of countable cofinality).
@@
This result is made possible by a technical advancement of the main crux of our previous paper (see [JL25, Theorem 4.11]). The major obstacle towards obtaining models in which a stationary set of a given cofinality is not in $I[\delta^{+}]$ is finding posets which can collapse successors of singular cardinals without making them $d$-approachable. Whether this was possible without preserving the successor of the singular as a cardinal had been open for around four decades and was a major obstacle both in non-approachable sets concentrating on specific cofinalities and iterating such constructions to obtain the failure of $\mathsf{AP}$ for many cardinals simultaneously. However, we solved that issue in [JL25] by constructing a poset which, under suitable hypotheses, collapses the successor of a singular cardinal of countable cofinality to any desired cofinality without making it $d$-approachable. That theorem relied on the tension between a weak form of the approximation property (see [JL25, Theorem 4.3.]) and our iteration not adding new short functions into small sets (see [JL25, Lemma 4.10]). In this work, we achieve a strengthening of [JL25, Theorem 4.3] (see Theorem 4.11) which allows us to dispense with [JL25, Lemma 4.10]. Using this, we are able to replace the Namba forcing used there by a new forcing notion which we call supercompact Namba forcing. This forcing is constructed using higher variants of the $\mathsf{LIP}$-ideals which we used in [JL25] and which have previously been considered by Matsubara (see [Mat02]). In technical terms, the poset from [JL25] uses infinitely many generically measurable cardinals while our new poset only uses a single generically supercompact cardinal (despite the latter being a stronger hypothesis it is much easier to obtain in our case).
@@ 1.1. Approachability, Colorings and Supercompact Cardinals
Let $\mu$ be a cardinal. A set $S\subseteq\mu^{+}$ is in the approachability ideal $I[\mu^{+}]$ if there is a club $C\subseteq\mu^{+}$ and a sequence $(x_{i})_{i<\mu^{+}}$ of elements of $\mu^{+}]^{<\mu}$ such that for any $\gamma\in S\cap C$, $\gamma$ is approachable with respect to $(x_{i})_{i<\mu^{+}}$, i.e. there is $A\subseteq\gamma$ unbounded of minimal ordertype such that whenever $\beta\in\gamma$, $A\cap\beta\in\{x_{i}\;|\;i<\gamma\}$.
@@
This ideal can be defined both for any cardinal $\mu$ but its behavior depends heavily on whether $\mu$ is regular or singular: If $\mu$ is regular, $E_{<\mu}^{\mu^{+}}\in I[\mu^{+}]$ and so $I[\mu^{+}]$ is completely determined by the membership of stationary subsets of $E_{\mu}^{\mu^{+}}$. On the other hand, if $\mu$ is singular, we can only prove that $E_{\leq\mu}^{\mu^{+}}\in I[\mu^{+}]$ which leaves open many more possibilities regarding the possible cofinalities for which $I[\mu^{+}]$ is nontrivial. As we will show, this result is sharp for successors of singular cardinals of countable cofinality.
@@
    $d(\alpha,\gamma)\leq\max\{d(\alpha,\beta),d(\beta,\gamma)\}$
@@
    $|\{\beta<\alpha\;|\;d(\beta,\alpha)\leq i\}|<\mu$
@@
    $\sup\{d(\beta,\alpha)\;|\;\beta\in A\cap\alpha\}<\operatorname{cf}(\mu)$
@@
Fact 1.3.
Let $\mu$ be a singular cardinal and $d\colon[\mu^{+}]^{2}\to\operatorname{cf}(\mu)$ a normal, subadditive coloring. An ordinal $\alpha<\mu^{+}$ is $d$-approachable if and only if $\operatorname{cf}(\alpha)\leq\operatorname{cf}(\mu)$ or there is a cofinal set $A\subseteq\alpha$ of minimal ordertype and $i<\operatorname{cf}(\mu)$ such that for all $\beta<\gamma$, both in $A$, $d(\beta,\gamma)\leq i$.
We now connect the previous two notions of approachability. In most of our models we will have that the $\mathsf{GCH}$ holds, so all our singular cardinals will be strong limits. This simplifies the previous concepts, because whenever $\mu$ is a singular strong limit, the following holds:
@@
    The approachability ideal $I[\mu^{+}]$ is equal to the weak approachability ideal $I[\mu^{+},\mu]$ (where instead of requiring $A\cap\beta\in\{x_{i}\;|\;i<\gamma\}$ for every $\beta<\gamma$ we merely require each $A\cap\beta$ to be a subset of $x_{i}$ for some $i<\gamma$ ).
@@
Let $\Theta$ and $x,y$ be given. Let $j\colon V\to M$ be a sufficiently supercompact embedding with critical point $\kappa$ such that $j(l)(\kappa)=y$. As before, the pointwise image $j[H^{V}(\Theta)]\prec H^{M}(j(\Theta))$ is a $j(\kappa)$-Magidor model. One shows by induction that $\pi_{j[H^{V}(\Theta)]}(j(x))=x$ whenever $x\in H^{V}(\Theta)$, which shows the statement (for more details, see [Jak25, Lemma 5.4.14]) ∎
@@
    $|\{\beta\in A\cap\alpha\;|\;d(\beta,\alpha)\leq i\}|<\mho$
@@
For (1), we first note that whenever $\mu\in M$ is a cardinal, $\operatorname{otp}(M\cap\mu)$ is a cardinal as well: Otherwise, there is $\alpha<\operatorname{otp}(M\cap\mu)$ and a well-order $X\subseteq\alpha\times\alpha$ with ordertype $\operatorname{otp}(M\cap\mu)$. It follows that $\alpha=\operatorname{otp}(M\cap\nu)$ for some $\nu<\mu$, $\nu\in M$ and there exists $B\subseteq M\cap\nu\times M\cap\nu$ with ordertype $\operatorname{otp}(M\cap\mu)$. Let $A\in M$ be such that $A\cap M=B$. Then by elementarity $A\subseteq\nu\times\nu$ and $\operatorname{otp}(M)=\mu$. But this contradicts the fact that $\mu$ is a cardinal.
@@
Lastly, we prove that Magidor models have transitive collapses which reflect many set-theoretical statements (this mirrors the fact that for all “bounded” statements $\phi$, one can find a sufficiently supercompact embedding $j\colon V\to M$ such that $V\models\phi\leftrightarrow M\models\phi$ due to the closure of $M$):
@@ 1.2. Forcing
We will also make use of projection analyses:
@@
The existence of a projection from $\mathbb{P}$ onto $\mathbb{Q}$ implies that any forcing extension by $\mathbb{Q}$ can again be extended by forcing to a forcing extension by $\mathbb{P}$. This is done using the quotient forcing:
@@
It follows that every forcing extension by $\mathbb{Q}$ can be once again forcing extended to a forcing extension by $\mathbb{P}$. This allows us to obtain regularity properties for a poset $\mathbb{Q}$ by finding a well-behaved poset $\mathbb{P}$ and a projection $\pi\colon\mathbb{P}\to\mathbb{Q}$. It is clear that in this situation, if $\mathbb{P}$ does not collapse certain cardinals or add sequences of ordinals of a certain length, neither does $\mathbb{Q}$, since any forcing extension by $\mathbb{Q}$ is contained in a forcing extension by $\mathbb{P}$.
@@
Let $\mathbb{P}:=(\mathbb{P},\leq_{\mathbb{P}})$ be the inverse limit of $(\mathbb{P}_{\alpha},\dot{\mathbb{Q}}_{\alpha})_{\alpha<\kappa}$. Denote by $\leq_{\mathbb{T}}$ the ordering on $\mathbb{T}$. It is clear that the identity is a function from $\mathbb{T}$ onto $\mathbb{P}$ and preserves the ordering. Let $p\in\mathbb{P}$ and let $t\leq_{\mathbb{P}}p$. We want to find $p^{\prime}\leq_{\mathbb{T}}p$ with $p^{\prime}\leq_{\mathbb{P}}t$. By induction on $\alpha$, let $p^{\prime}(\alpha)$ be a $\mathbb{P}_{\alpha}$-name for an element of $\dot{\mathbb{Q}}_{\alpha}$ forced by $p^{\prime}\upharpoonright\alpha$ to be equal to $t(\alpha)$ and by conditions incompatible with $p^{\prime}\upharpoonright\alpha$ to be equal to $p(\alpha)$. Then clearly $p^{\prime}\leq_{\mathbb{T}}p$ (since $p^{\prime}(\alpha)$ is in any case forced to be below $p(\alpha)$) and $p^{\prime}\leq_{\mathbb{P}}t$ (since $p^{\prime}(\alpha)$ is always forced by $p^{\prime}\upharpoonright\alpha$ to be equal to $t(\alpha)$). ∎
We use the following new closure property simply because it fits our proofs:
@@
It is easy to see that strong directed closure is productive. Additionally, the usual proof of the iterability of directed closure goes through for strong directed closure as well:
@@
Another proof which adapts to this new definition is that strong directed closure is inherited by the termspace forcing:

### v2

@@ Abstract.
The notion of independence is fundamental to set-theoretic research. It refers to the phenomenon that many set-theoretic statements are neither proven nor refuted by $\mathsf{ZFC}$, the most commonly accepted axioms of set theory. One particular theme in the study of independence statements is the fact that successors of singular cardinals are often much less malleable than successors of regular cardinals – recall that a cardinal $\kappa$ is singular if we can write $\kappa=\bigcup_{i\in x}y_{i}$, where $x$ and every $y_{i}$ has size $\kappa$ and regular otherwise. Additionally, many independence results for regular cardinals can not be obtained for singulars. The most famous example of this phenomenon is the combination of two results due to William Easton and Jack Silver: Easton showed that the continuum function – mapping a cardinal $\kappa$ to the size of its powerset – when restricted to the class of regular cardinals can be equal to any function, provided that function is nondecreasing and maps every cardinal to a cardinal with a larger cofinality. On the other hand, Silver showed that whenever $\kappa$ is a singular cardinal with uncountable cofinality – meaning that any function from $\omega$ into $\kappa$ is bounded – and $2^{\nu}=\nu^{+}$ for every cardinal $\nu<\kappa$, then $2^{\kappa}=\kappa^{+}$. Thus, Easton’s result cannot be replicated for the class of all cardinals. And even at $\aleph_{\omega}$, the first singular cardinal, Saharon Shelah demonstrated provable constraints regarding the continuum function: He showed that, assuming $2^{\aleph_{n}}<\aleph_{\omega}$ for all $n<\omega$, $2^{\aleph_{\omega}}<\aleph_{\omega_{4}}$.
@@
Another example highlighting the difference between successors of singular cardinals and successors of regular cardinals is Shelah’s approachability ideal: Shelah introduced the ideal $I[\kappa^{+}]$ in [She79] in order to study the indestructibility of stationary subsets of $\kappa^{+}$ under sufficiently closed forcings and did so both for singular and regular cardinals $\kappa$. Despite the definition of $I[\kappa^{+}]$ being the same in both cases, its behavior differs greatly depending on whether $\kappa$ is regular or singular: When $\kappa$ is a regular cardinal, the set $E_{<\kappa}^{\kappa^{+}}$, consisting of all ordinals below $\kappa^{+}$ with cofinality $\kappa$, is always a member of $I[\kappa^{+}]$ and so the behavior of $I[\kappa^{+}]$ is determined by which subsets of $E_{\kappa}^{\kappa^{+}}$ it contains. On the other hand, if $\kappa$ is singular, it was only proven that the set $E_{\leq\operatorname{cf}(\kappa)}^{\kappa^{+}}$ belongs to $I[\kappa^{+}]$ and so many more options regarding the behavior of $I[\kappa^{+}]$ seem possible.
@@
Recently, in joint work of Maxwell Levine and the author (see [JL25]), we solved an almost four-decade-old open problem due to Saharon Shelah by constructing a model – for an arbitrary $n\in(0,\omega)$ – where $E_{\aleph_{n}}^{\aleph_{\omega+1}}\notin I[\aleph_{\omega+1}]$. Previously it had only been known that there could be stationarily many non-approachable points of cofinality $\aleph_{1}$ in $\aleph_{\omega+1}$. Our result begs the natural question (which was also raised by Mitchell in [Mit09] and by Dobrinen in private communication) whether it is consistent for a singular cardinal $\kappa$ of uncountable cofinality that $E_{\mu}^{\kappa^{+}}\notin I[\kappa^{+}]$ for every regular cardinal $\mu\in(\aleph_{0},\kappa)$.
@@
In particular, whenever $\delta$ is singular with countable cofinality, $\mathsf{AP}_{\delta}$, which states that $\delta^{+}\notin I[\delta^{+}]$, does not hold. Since $\mathsf{AP}_{\delta}$ is implied by the weak square principle $\square_{\delta}^{*}$, the use of large cardinals is provably necessary for this result (see e.g. [Sar14]).
Our main result is made possible by a technical advancement of the main crux of our previous paper (see [JL25, Theorem 4.11]). The major obstacle towards obtaining models in which a stationary set of a given cofinality is not in $I[\delta^{+}]$ is finding posets which can collapse successors of singular cardinals without making them $d$-approachable. Whether this was possible without preserving the successor of the singular as a cardinal had been open for around four decades and was a major obstacle both in non-approachable sets concentrating on specific cofinalities and iterating such constructions to obtain the failure of $\mathsf{AP}$ for many cardinals simultaneously. However, we solved that issue in [JL25] by constructing a poset which, under suitable hypotheses, collapses the successor of a singular cardinal of countable cofinality to any desired cofinality without making it $d$-approachable. That theorem relied on the tension between a weak form of the approximation property (see [JL25, Theorem 4.3.]) and our iteration not adding new short functions into small sets (see [JL25, Lemma 4.10]). In this work, we achieve a strengthening of [JL25, Theorem 4.3] (see Theorem 4.12) which allows us to dispense with [JL25, Lemma 4.10]. Using this, we are able to replace the Namba forcing used there by a new forcing notion which, similarly to supercompact Prikry forcing, uses elements of $\lambda]^{<\kappa}$ as opposed to ordinals. This forcing is constructed using higher variants of the $\mathsf{LIP}$-ideals which we used in [JL25] and which have previously been considered by Matsubara (see [Mat02]). In technical terms, the poset from [JL25] uses infinitely many generically measurable cardinals while our new poset only uses a single generically supercompact cardinal. Despite the latter being a stronger hypothesis in terms of its consistency strength, it is much easier to obtain in our case.
@@
Acknowledgments.
The author wants to thank the referee for their diligent reading and valuable comments, leading to an improvement of the manuscript.
@@ 1.1. Approachability, Colorings and Supercompact Cardinals
Let $\mu$ be a cardinal. A set $S\subseteq\mu^{+}$ is in the approachability ideal $I[\mu^{+}]$ if there is a club $C\subseteq\mu^{+}$ and a sequence $(x_{i})_{i<\mu^{+}}$ of elements of $\mu^{+}]^{<\mu}$ such that for any $\gamma\in S\cap C$, $\gamma$ is approachable with respect to $(x_{i})_{i<\mu^{+}}$, i.e. there is $A\subseteq\gamma$ unbounded of minimal order-type such that whenever $\beta\in\gamma$, $A\cap\beta\in\{x_{i}\;|\;i<\gamma\}$.
@@
Although not directly relevant to the contents of this manuscript, the approachability ideal $I[\mu^{+}]$ for singular $\mu$ is intimately related to the concept of scales:
Definition 1.2.
Let $\mu$ be a singular cardinal with countable cofinality. Let $(\mu_{n})_{n\in\omega}$ be an increasing sequence of regular cardinals converging to $\mu$. A $(\mu^{+},(\mu_{n})_{n\in\omega})$-scale is a sequence $(f_{\alpha})_{\alpha<\mu^{+}}$ such that:
  1. (1)
    For all $\alpha<\mu^{+}$, $f_{\alpha}\in\prod_{n\in\omega}\mu_{n}$;
  2. (2)
    For all $\alpha<\beta<\mu^{+}$, $f_{\alpha}<^{*}f_{\beta}$, i.e. there is $k\in\omega$ such that $f_{\alpha}(n)<f_{\beta}(n)$ for all $n\geq k$;
  3. (3)
    For all $g\in\prod_{n\in\omega}\mu_{n}$ there is $\alpha<\mu^{+}$ such that $g<^{*}f_{\alpha}$.
Given a $(\mu^{+},(\mu_{n})_{n\in\omega})$-scale $(f_{\alpha})_{\alpha<\mu^{+}}$, we say that an ordinal $\gamma<\mu^{+}$ is good for $(f_{\alpha})_{\alpha<\mu^{+}}$ if there exists an unbounded $A\subseteq\gamma$ and $k\in\omega$ such that for all $n\geq k$, the sequence $(f_{\alpha}(n))_{\alpha\in A}$ is strictly increasing. $\gamma<\mu^{+}$ is bad for $(f_{\alpha})_{\alpha<\mu^{+}}$ if it is not good for $(f_{\alpha})_{\alpha<\mu^{+}}$.
Given a $(\mu^{+},(\mu_{n})_{n\in\omega})$-scale $\vec{f}$ and a set $S\subseteq\mu^{+}$ in $I[\mu^{+}]$, Shelah showed that there is a club subset $C\subseteq\mu^{+}$ such that every $\gamma\in S\cap C$ is good for $\vec{f}$. On the other hand, answering an old question due to Shelah, Maxwell Levine and the author showed in [JL25] that it is consistent to have a $(\aleph_{\omega+1},(\aleph_{n})_{n\in\omega})$-scale $\vec{f}$ such that club many points in $\aleph_{\omega+1}$ are good for $\vec{f}$, but $\aleph_{\omega+1}\notin I[\aleph_{\omega+1}]$. In our model, there will also be many instances of this phenomenon: By a folklore result111see https://mathoverflow.net/q/296225, whenever $\vec{f}$ is a $(\mu^{+},(\mu_{n})_{n\in\omega})$-scale, there is a club subset $C\subseteq\mu^{+}$ such that any $\gamma\in C\cap E_{\geq\aleph_{3}}^{\mu^{+}}$ is good for $\vec{f}$. Despite this, in our model, this set will clearly be outside of $I[\mu^{+}]$.
The approachability ideal can be defined for any cardinal $\mu$ but its behavior depends heavily on whether $\mu$ is regular or singular: If $\mu$ is regular, $E_{<\mu}^{\mu^{+}}\in I[\mu^{+}]$ (see [She91, Lemma 4.4]) and so $I[\mu^{+}]$ is completely determined by the membership of stationary subsets of $E_{\mu}^{\mu^{+}}$. On the other hand, if $\mu$ is singular, the only known results are that $E_{\leq\operatorname{cf}(\mu)}^{\mu^{+}}$ is an element of $I[\mu^{+}]$ and for any regular uncountable cardinal $\delta<\mu$, $I[\mu^{+}]$ contains a stationary subset of $E_{\delta}^{\mu^{+}}$ (see [Cum05, Theorem 9.2]). As we will show, this result is sharp for successors of singular cardinals of countable cofinality.
@@
    $d(\alpha,\gamma)\leq\max\{d(\alpha,\beta),d(\beta,\gamma)\};$
@@
    $|\{\beta<\alpha\;|\;d(\beta,\alpha)\leq i\}|<\mu;$
@@
    $\sup\{d(\beta,\alpha)\;|\;\beta\in A\cap\alpha\}<\operatorname{cf}(\mu);$
@@
Fact 1.4.
Let $\mu$ be a singular cardinal and $d\colon[\mu^{+}]^{2}\to\operatorname{cf}(\mu)$ a normal, subadditive coloring. An ordinal $\alpha<\mu^{+}$ is $d$-approachable if and only if $\operatorname{cf}(\alpha)\leq\operatorname{cf}(\mu)$ or there is a cofinal set $A\subseteq\alpha$ of minimal order-type and $i<\operatorname{cf}(\mu)$ such that for all $\beta<\gamma$, both in $A$, $d(\beta,\gamma)\leq i$.
In the case where the singular cardinal $\mu$ under consideration is a strong limit, the previous concepts are connected as follows:
@@
    The approachability ideal $I[\mu^{+}]$ is equal to the weak approachability ideal $I[\mu^{+},\mu]$. This ideal is defined almost the same way as $I[\mu^{+}]$ but instead of requiring $A\cap\beta\in\{x_{i}\;|\;i<\gamma\}$ for every $\beta<\gamma$ we merely require each $A\cap\beta$ to be a subset of $x_{i}$ for some $i<\gamma$.
@@
Let $\Theta$ and $x,y$ be given. Let $j\colon V\to M$ be a sufficiently supercompact embedding with critical point $\kappa$ such that $j(l)(\kappa)=y$. As before, the pointwise image $j[H^{V}(\Theta)]\prec H^{M}(j(\Theta))$ is a $j(\kappa)$-Magidor model. One shows by induction that $\pi_{j[H^{V}(\Theta)]}(j(a))=a$ whenever $a\in H^{V}(\Theta)$, which shows the statement (for more details, see [Jak25, Lemma 5.4.14]) ∎
@@
    $|\{\beta\in A\cap\alpha\;|\;d(\beta,\alpha)\leq i\}|<\mho.$
@@
For (1), we first note that whenever $\mu\in M$ is a cardinal, $\operatorname{otp}(M\cap\mu)$ is a cardinal as well: Otherwise, there is $\alpha<\operatorname{otp}(M\cap\mu)$ and a well-order $X\subseteq\alpha\times\alpha$ with order-type $\operatorname{otp}(M\cap\mu)$. It follows that $\alpha=\operatorname{otp}(M\cap\nu)$ for some $\nu<\mu$, $\nu\in M$ and there exists $B\subseteq M\cap\nu\times M\cap\nu$ with order-type $\operatorname{otp}(M\cap\mu)$. Let $A\in M$ be such that $A\cap M=B$. Then by elementarity $A\subseteq\nu\times\nu$ and $\operatorname{otp}(M)=\mu$. But this contradicts the fact that $\mu$ is a cardinal.
@@
Lastly, we prove that Magidor models have transitive collapses which reflect many set-theoretical statements. This mirrors the fact that for all “bounded” statements $\phi$, one can find a sufficiently supercompact embedding $j\colon V\to M$ such that $V\models\phi\leftrightarrow M\models\phi$ due to the closure of $M$:
@@ 1.2. Forcing
We will also make use of projection analyses. In the literature, there are two common ways of defining projections, one due to Abraham (see [AM10, Page 335]) and a slightly more general one due to Cummings (see [Cum10, Definition 5.2]). In this work, we will use the definition due to Cummings, since the results it provides are sufficient for us. [^2] See https://mathoverflow.net/q/466836 for a discussion regarding the discrepancy.
@@
The existence of a projection from $\mathbb{P}$ onto $\mathbb{Q}$ implies that, after forcing with $\mathbb{Q}$, the resulting model can once again be extended by forcing to a model which is a forcing extension by $\mathbb{P}$. This is done using the quotient forcing:
@@
It follows that, after forcing with $\mathbb{Q}$, we can perform further forcing to reach a forcing extension by $\mathbb{P}$. This allows us to obtain regularity properties for a poset $\mathbb{Q}$ by finding a well-behaved poset $\mathbb{P}$ and a projection $\pi\colon\mathbb{P}\to\mathbb{Q}$. It is clear that in this situation, if $\mathbb{P}$ does not collapse certain cardinals or add sequences of ordinals of a certain length, neither does $\mathbb{Q}$, since any forcing extension by $\mathbb{Q}$ is contained in a forcing extension by $\mathbb{P}$.
@@
Let $\mathbb{P}:=(\mathbb{P},\leq_{\mathbb{P}})$ be the inverse limit of $(\mathbb{P}_{\alpha},\dot{\mathbb{Q}}_{\alpha})_{\alpha<\kappa}$. Denote by $\leq_{\mathbb{T}}$ the ordering on $\mathbb{T}$. It is clear that the identity is a function from $\mathbb{T}$ onto $\mathbb{P}$ and preserves the ordering. Let $p\in\mathbb{P}$ and let $t\leq_{\mathbb{P}}p$. We want to find $p^{\prime}\leq_{\mathbb{T}}p$ with $p^{\prime}\leq_{\mathbb{P}}t$. By induction on $\alpha$, let $p^{\prime}(\alpha)$ be a $\mathbb{P}_{\alpha}$-name for an element of $\dot{\mathbb{Q}}_{\alpha}$ forced by $p^{\prime}\upharpoonright\alpha$ to be equal to $t(\alpha)$ and by conditions incompatible with $p^{\prime}\upharpoonright\alpha$ to be equal to $p(\alpha)$. Since $p^{\prime}(\alpha)$ is in any case forced to be below $p(\alpha)$, $p^{\prime}\leq_{\mathbb{T}}p$. Since $p^{\prime}(\alpha)$ is always forced by $p^{\prime}\upharpoonright\alpha$ to be equal to $t(\alpha)$, $p^{\prime}\leq_{\mathbb{P}}t$. ∎
The proof of the main theorem is done by successively forcing, for a regular cardinal $\mu$, that for every singular $\delta>\mu$ the set $E_{\mu}^{\delta^{+}}$ is not in $I[\delta^{+}]$. In order to prove that the desired conclusion holds, we must therefore show that the tail of the iteration does not destroy the stationarity of the already added sets that lie outside of $I[\delta^{+}]$. We will do so by proving that a certain elementary embedding lifts through the forcing extension, which requires us to show that the tail forcing is sufficiently directed-closed:
Definition 1.22.
Let $\mathbb{P}$ be a poset. $X\subseteq\mathbb{P}$ is directed if whenever $p,q\in X$, there is $r\in X$ such that $r\leq p,q$. If $\mu$ is a cardinal, $\mathbb{P}$ is $\mu-$directed closed if whenever $X\subseteq\mathbb{P}$ is directed and $|X|<\mu$, there exists $p\in\mathbb{P}$ such that $p\leq q$ for every $q\in X$.
In the proof of Lemma 3.3, there is a single point where the established notion of directed closure seems to be insufficient, unless the poset under consideration is separative. Since the closure of a poset might fail to be inherited by its separative quotient333see https://mathoverflow.net/q/11505, we are defining the following strengthening of directed closure instead and using it to prove that the separative quotient of our considered poset is directed-closed.
@@
It is easy to see that strong directed closure is productive. Additionally, the usual proof of the iterability of directed closure goes through for strong directed closure as well. Note that, in general, this does not automatically follow from the iterability of directed closure, since two-step iterations typically fail to be separative.
@@
Another proof which adapts to this new definition is that strong directed closure is inherited by the termspace forcing. Note that, again, this is not automatic, since in general the termspace forcing is not separative.